OpenStudy (anonymous):

Simplify the following expression. 7/100 dived 7/12

OpenStudy (anonymous):

7/100 : 7/12 = 7/100 * 12/7 = 1/100 * 12/1 = 1/25 * 3/1 = 3/25

OpenStudy (anonymous):

= 0.00083333333